Find remainder when 55*347*90divided by 7 without multiplying these numbers
Answer by Edwin McCravy(20065)

55*347*90 = (49+6)(343+4)(84+6) = (7*7+6)(49*7+4)(12*7+6)

When that is multiplied out, all terms will be divisible by 7 except
the last term 6*4*6 = 144, which leaves remainder 4 when divided by 7.

Answer: 4

Edwin